# Quantum Pseudorandomness and Classical Complexity

### v4

@@ 1.1 Main results
Theorem 2 (Informal version of Theorem 29 and Theorem 32).
@@
    $\mathsf{BQP}^{\mathcal{O}}=\mathsf{QMA}^{\mathcal{O}}$, [^3] In fact, we also have $\mathsf{PromiseBQP}^{\mathcal{O}}=\mathsf{PromiseQMA}^{\mathcal{O}}$, which is stronger. For the purposes of this exposition, however, we restrict attention to classes of languages. and
@@
Theorem 2 illustrates a stark contrast between quantum and classical cryptography, because the existence of hard problems in $\mathsf{NP}$ is necessary to have classical cryptography that is secure against polynomial-time adversaries. One can view our result as evidence that the same is not necessary for the existence of quantum cryptography; perhaps weaker assumptions suffice. Indeed, because a major goal in cryptography is to build cryptosystems from minimal computational assumptions, Theorem 2 has served as the primary motivation for many recent works that have built cryptography from pseudorandom states and unitaries [AQY22, MY22b, BCQ23, MY22a, HMY22]. Note that these works all appeared after this work was originally published [Kre21a].
@@ 1.2 Application: hyperefficient shadow tomography
Nevertheless, our results imply that black-box hyperefficient shadow tomography is impossible even in this setting where we have state preparation access to the observables. This follows from the simple observation that hyperefficient shadow tomography of this form would suffice to break PRS ensembles with a $\mathsf{QCMA}$ oracle.
@@ 1.3 Our techniques
Approximate $t$-designs
Approximate $t$-designs play an important role in the proofs of both Theorem 1 and Theorem 2. So, in Section 3, we give formal definitions of $t$-designs and prove some of their useful properties in the context of quantum query complexity. In particular, we establish conditions under which which substituting the Haar measure with a $t$-design yields a relative-error approximation to the acceptance probability of a quantum query algorithm. Several authors have implicitly assumed without proof that this property holds, e.g. [BHH16b, AMR20], and also an earlier version of this work [Kre21a]. We consider it valuable to place these results on more rigorous footing, and believe that the results about $t$-designs proved herein could find independent uses in other complexity-theoretic contexts.
Breaking pseudorandomness with $\mathsf{PP}$
@@
Fix $M$ different observables $O_{1},O_{2},\ldots,O_{M}$ and an unknown $n$-qubit mixed state $\rho$. Then there exists a quantum algorithm that performs $T=O(\log(M/\delta)/\varepsilon^{2}\cdot\max_{i}\mathrm{Tr}(O_{i}^{2}))$ single-copy measurements in random Clifford bases [^5] Recall that the Clifford group is the group of unitary transformation generated by Hadamard, phase, and CNOT gates. A Clifford basis is any basis that can be obtained from the computational basis via multiplication by an element of the Clifford group. of $\rho$, and uses the measurement results (called classical shadows) to estimate the quantities $\mathrm{Tr}(O_{1}\rho),\mathrm{Tr}(O_{2}\rho),\ldots,\mathrm{Tr}(O_{M}\rho)$, such that with probability at least $1-\delta$, all of the $M$ quantities are correct up to additive error $\varepsilon$.
If $\{|\varphi_{k}\rangle\}_{k\in\{0,1\}^{\kappa}}$ is a pseudorandom state ensemble, then by choosing $O_{k}=|\varphi_{k}\rangle\langle\varphi_{k}|$ for each $k\in\{0,1\}^{\kappa}$ to be the list of observables, we can use the above algorithm to determine whether $\rho$ is close to one of the states in the PRS ensemble. A Haar-random state will be far from all of the pseudorandom states with overwhelming probability. Hence, Theorem 4 implies the existence of an algorithm that distinguishes the pseudorandom and Haar-random ensembles, by performing a polynomial number of random Clifford measurements and analyzing the results. The key observation is that the Clifford measurements can be performed efficiently, even though the resulting analysis (which operates on purely classical information) might be computationally expensive.
@@
Instantiating pseudorandomness with $\mathsf{BQP}=\mathsf{QMA}$
@@
For our second result (Theorem 2), the oracle construction we use is simple to describe. The oracle $\mathcal{O}$ consists of two parts: a quantum oracle $\mathcal{U}=\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$, where each $\mathcal{U}_{n}$ consists of $2^{n}$ different Haar-random $n$-qubit unitary matrices, and a classical oracle $\mathcal{P}$ that is an arbitrary $\mathsf{PSPACE}$-complete language. We prove that Theorem 2 holds with probability $1$ over the choice of $\mathcal{U}$.
@@
Showing that PRUs exist relative to $(\mathcal{U},\mathcal{P})$ is reasonably straightforward. The proof uses the BBBV theorem (i.e. the optimality of Grover’s algorithm) [BBBV97], and is analogous to showing that one-way functions or pseudorandom generators exist relative to a random classical oracle, as was shown by Impagliazzo and Rudich [IR89]. We only rigorously prove security against adversaries with classical advice, though we believe that the framework of Chung, Guo, Liu, and Qian [CGLQ20] should yield a security proof against adversaries with quantum advice.
@@
Slightly more technically involved is proving that $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{P}}=\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{P}}$. To do so, we argue that a $\mathsf{QMA}$ verifier is not substantially more powerful than a $\mathsf{BQP}$ machine at learning nontrivial properties of $\mathcal{U}$. More precisely, we argue that if a $\mathsf{QMA}$ verifier $\mathcal{V}$ makes $T$ queries to $\mathcal{U}_{n}$ for some $n\in\mathbb{N}$, then either (1) $n=O(\log T)$ is sufficiently small that $\mathrm{poly}(T)$ queries to $\mathcal{U}_{n}$ actually suffice to learn $\mathcal{U}_{n}$ to inverse-polynomial precision, or else (2) $n=\omega(\log t)$ is sufficiently large that with high probability, the maximum acceptance probability of $\mathcal{V}$ (over the choice of Merlin’s witness) is close to the average maximum acceptance probability of $\mathcal{V}$ when $\mathcal{U}_{n}$ is replaced by a random set of matrices sampled from the Haar measure. We prove this as a consequence of the extremely strong concentration of measure properties exhibited by the Haar measure [Mec19].
@@
This allows a $\mathsf{BQP}^{\mathcal{U},\mathcal{P}}$ machine to approximate the maximum acceptance probability of $\mathcal{V}^{\mathcal{U},\mathcal{P}}$ as follows. In case (1), the $\mathsf{BQP}^{\mathcal{U},\mathcal{P}}$ machine first queries $\mathcal{U}_{n}$ enough times to learn a unitary transformation $\widetilde{\mathcal{U}}_{n}$ that is close to $\mathcal{U}_{n}$, and then hard codes $\widetilde{\mathcal{U}}_{n}$ into a new $\mathsf{QMA}^{\mathcal{P}}$ verifier $\widetilde{\mathcal{V}}$ that simulates $\mathcal{V}$ by replacing queries to $\mathcal{U}_{n}$ with calls to $\widetilde{\mathcal{U}_{n}}$. In case (2), the $\mathsf{BQP}^{\mathcal{U},\mathcal{P}}$ machine similarly constructs a new $\mathsf{QMA}^{\mathcal{P}}$ verifier $\widetilde{\mathcal{V}}$, instead simulating $\mathcal{V}$ by replacing queries to $\mathcal{U}_{n}$ with unitaries chosen from an approximate polynomial design. [^6] Technically, this requires choosing a random element of the polynomial design for each $x\in\{0,1\}^{n}$ by means of a random oracle, so we use Zhandry’s strategy [Zha12] to simulate $T$ quantum queries to a random oracle using a $2T$-wise independent function. In both cases, $\widetilde{\mathcal{V}}$ defines a $\mathsf{QMA}^{\mathcal{P}}$ problem. Because $\mathcal{P}$ is $\mathsf{PSPACE}$-complete, $\mathsf{BQP}^{\mathcal{P}}=\mathsf{QMA}^{\mathcal{P}}=\mathsf{PSPACE}$, and therefore this problem can be decided with a single query to $\mathcal{P}$.
The astute reader may notice that this proof works for more general choices of $\mathcal{P}$: it shows that for any oracle $\mathcal{P}$, if $\mathsf{BQP}^{\mathcal{P}}=\mathsf{QMA}^{\mathcal{P}}$, then $\mathsf{BQP}^{\mathcal{\mathcal{U},P}}=\mathsf{QMA}^{\mathcal{\mathcal{U},P}}$ with probability $1$ over the choice of $\mathcal{U}$. An interesting consequence is the special case when $\mathcal{P}$ is trivial.
Corollary 5.
If $\mathsf{BQP}=\mathsf{QMA}$, then $\mathsf{BQP}^{\mathcal{U}}=\mathsf{QMA}^{\mathcal{\mathcal{U}}}$ with probability $1$ over the choice of $\mathcal{U}$.
In words, if $\mathsf{BQP}=\mathsf{QMA}$ in the unrelativized world, then the complexity classes also coincide relative to a collection $\mathcal{U}$ of Haar-random oracles. Or, viewed another way, separating $\mathsf{BQP}$ from $\mathsf{QMA}$ relative to $\mathcal{U}$ requires separating them in the unrelativized world. This is in stark contrast to the case of random classical oracles, where we can prove unconditionally that for a uniformly random oracle $\mathcal{O}$, $\mathsf{BQP}^{\mathcal{O}}\neq\mathsf{QMA}^{\mathcal{O}}$ (and indeed, $\mathsf{NP}^{\mathcal{O}}\not\subset\mathsf{BQP}^{\mathcal{O}}$) with probability $1$ over $\mathcal{O}$ [BBBV97].
@@ 1.4 Open problems
Can we prove a similar result to Theorem 2 using a classical oracle, for either PRUs or PRSs? Attempting to resolve this question seems to run into many of the same difficulties that arise in constructing a classical oracle separation between $\mathsf{QCMA}$ and $\mathsf{QMA}$, which also remains an open problem [AK07]. For one, as pointed out in [AK07], we do not even know whether every $n$-qubit unitary transformation can be approximately implemented in $\mathrm{poly}(n)$ time relative to some classical oracle—this is sometimes known as the unitary synthesis problem [Aar16, Ros21]. Even if one could resolve this, it is not clear whether the resulting PRUs or PRSs would be secure against adversaries with the power of $\mathsf{QMA}$. For instance, we show in Appendix A that an existing construction of PRSs, whose security is provable in the random oracle model [BS19], can be broken with an $\mathsf{NP}$ oracle. Nevertheless, recent work by Kretschmer, Qian, Sinha, and Tal [KQST23] makes progress on this question by constructing an oracle relative to which $\mathsf{P}=\mathsf{NP}$ and a weaker version of pseudorandom states exist with only single-copy security.
What else can be said about the hardness of learning quantum states and unitary transformations, either in the worst case or on average? A related question is to explore the hardness of problems involving quantum meta-complexity: that is, problems that themselves encode computational complexity or difficulty. Consider, for example, a version of the minimum circuit size problem ($\mathsf{MCSP}$) for quantum states: given copies of a pure quantum state $|\psi\rangle$, determine the size of the smallest quantum circuit that approximately outputs $|\psi\rangle$. If PRSs exist, then this task should be hard, but placing an upper bound on the complexity of this task might be difficult in light of our results. We view this problem as particularly intriguing because it does not appear to have an obvious classical analogue, and also because of its relevance to the wormhole growth paradox and Susskind’s Complexity=Volume conjecture in AdS/CFT [BFV20, Sus16b, Sus16a]. A number of recent breakthroughs in complexity theory have involved ideas from meta-complexity (see surveys by Allender [All17, All20]), and it would be interesting to see which of these techniques could be ported to the quantum setting.
@@ 2.2 Quantum information
We let $\mathrm{TD}(\rho,\sigma)$ denote the trace distance between density matrices $\rho$ and $\sigma$. We use $||M||_{F}=\sqrt{\mathrm{Tr}\left(M^{\dagger}M\right)}$ to denote the Frobenius norm of a matrix $M$.
@@
$\left|\left|\mathcal{A}\right|\right|_{\diamond}=\sup_{\mathrm{Tr}(\rho)=1,\rho\succeq 0}(\mathcal{A}\otimes I)(\rho),$
@@
$\mathrm{TD}(\mathcal{A}(\rho),\mathcal{B}(\rho))\leq\left|\left|\mathcal{A}-\mathcal{B}\right|\right|_{\diamond}.$
@@
A consequence of 7 is the following well-known bound relating distance in diamond norm to the Frobenius norm for unitary channels. We provide a proof for completeness.
@@
Let $d$ be the distance in the complex plane between 00 and the polygon whose vertices are $\lambda_{1},\ldots,\lambda_{N}$. Then from 7 we may conclude:

### v5

@@ 1.1 Main results
Theorem 2 (Informal version of Theorems 30 and 33).
@@
    $\mathsf{BQP}^{\mathcal{O}}=\mathsf{QMA}^{\mathcal{O}}$, and
@@
In fact, our oracle $\mathcal{O}$ also satisfies $\mathsf{PromiseBQP}^{\mathcal{O}}=\mathsf{PromiseQMA}^{\mathcal{O}}$, which is stronger. For the sake of clarity, in this introduction we will only state our results in terms of classes of languages (e.g. $\mathsf{QMA}$) instead of classes of promise problems (e.g. $\mathsf{PromiseQMA}$), unless the distinction matters.
@@
Theorem 2 illustrates a stark contrast between quantum and classical cryptography, because the existence of hard problems in $\mathsf{NP}$ is necessary to have classical cryptography that is secure against polynomial-time adversaries. One can view our result as evidence that the same is not necessary for the existence of quantum cryptography; perhaps weaker assumptions suffice. Indeed, because a major goal in cryptography is to build cryptosystems from minimal computational assumptions, Theorem 2 has served as the primary motivation for many recent works that have built cryptography from pseudorandom states and unitaries [AQY22, MY22b, BCQ23, MY22a, HMY23]. Note that these works all appeared after this work was originally published [Kre21a].
@@ 1.2 Application: hyperefficient shadow tomography
Nevertheless, our results imply that black-box hyperefficient shadow tomography is impossible even in this setting where we have state preparation access to the observables. This follows from the simple observation that hyperefficient shadow tomography of this form would suffice to break PRS ensembles with a ($\mathsf{Promise}$)$\mathsf{QCMA}$ oracle.
@@ 1.3 Our techniques
1.3.1 Approximate $t$-designs
Approximate $t$-designs play a role in the proof of Theorem 1. So, in Section 3, we give formal definitions of $t$-designs and prove some of their useful properties in the context of quantum query complexity. In particular, we establish conditions under which which substituting the Haar measure with a $t$-design yields a relative-error approximation to the acceptance probability of a quantum query algorithm. Several authors have implicitly assumed without proof that this property holds, e.g. [BHH16b, AMR20], and also an earlier version of this work [Kre21a]. We consider it valuable to place these results on more rigorous footing, and believe that the results about $t$-designs proved herein could find independent uses in other complexity-theoretic contexts.
1.3.2 Breaking pseudorandomness with $\mathsf{PP}$
@@
Fix $M$ different observables $O_{1},O_{2},\ldots,O_{M}$ and an unknown $n$-qubit mixed state $\rho$. Then there exists a quantum algorithm that performs $T=O(\log(M/\delta)/\varepsilon^{2}\cdot\max_{i}\mathrm{Tr}(O_{i}^{2}))$ single-copy measurements in random Clifford bases [^4] Recall that the Clifford group is the group of unitary transformations generated by Hadamard, phase, and CNOT gates. A Clifford basis is any basis that can be obtained from the computational basis via multiplication by an element of the Clifford group. of $\rho$, and uses the measurement results (called classical shadows) to estimate the quantities $\mathrm{Tr}(O_{1}\rho),\mathrm{Tr}(O_{2}\rho),\ldots,\mathrm{Tr}(O_{M}\rho)$, such that with probability at least $1-\delta$, all of the $M$ quantities are correct up to additive error $\varepsilon$.
If $\{|\varphi_{k}\rangle\}_{k\in\{0,1\}^{\kappa}}$ is a pseudorandom state ensemble, then by choosing $O_{k}=|\varphi_{k}\rangle\langle\varphi_{k}|$ for each key $k\in\{0,1\}^{\kappa}$ to be the list of observables, we can use the above algorithm to determine whether $\rho$ is close to one of the states in the PRS ensemble. A Haar-random state will be far from all of the pseudorandom states with overwhelming probability. Hence, Theorem 4 implies the existence of an algorithm that distinguishes the pseudorandom and Haar-random ensembles, by performing a polynomial number of random Clifford measurements and analyzing the results. The key observation is that the Clifford measurements can be performed efficiently, even though the resulting analysis (which operates on purely classical information) might be computationally expensive.
@@
1.3.3 Instantiating pseudorandomness with $\mathsf{BQP}=\mathsf{QMA}$
@@
For our second result (Theorem 2), the oracle $\mathcal{O}$ that we construct consists of two parts: a quantum oracle $\mathcal{U}=\{\mathcal{U}_{n}\}_{n\in\mathbb{N}}$, where each $\mathcal{U}_{n}$ consists of $2^{n}$ different Haar-random $n$-qubit unitary matrices, and a classical oracle (i.e. a language) $\mathcal{C}$ that we build independently of $\mathcal{U}$. We prove that Theorem 2 holds with probability $1$ over the choice of $\mathcal{U}$.
@@
Showing that PRUs exist relative to $(\mathcal{U},\mathcal{C})$ is reasonably straightforward. Notably, the security proof does not depend on the choice of $\mathcal{C}$, so long as $\mathcal{C}$ is independent of the randomly sampled $\mathcal{U}$. The proof uses the BBBV theorem (i.e. the optimality of Grover’s algorithm) [BBBV97], and is analogous to showing that one-way functions or pseudorandom generators exist relative to a random classical oracle, as was shown by Impagliazzo and Rudich [IR89]. We only rigorously prove security against adversaries with classical advice, though we believe that the framework of Chung, Guo, Liu, and Qian [CGLQ20] should yield a security proof against adversaries with quantum advice.
@@
Slightly more technically involved is proving that $\mathsf{BQP}^{\mathcal{U},\mathcal{C}}=\mathsf{QMA}^{\mathcal{U},\mathcal{C}}$. To do so, we argue that a $\mathsf{QMA}$ verifier is not substantially more powerful than a $\mathsf{BQP}$ machine at learning nontrivial properties of $\mathcal{U}$. More precisely, we argue that if a $\mathsf{QMA}$ verifier $\mathcal{V}$ makes $T$ queries to $\mathcal{U}_{n}$ for some $n\in\mathbb{N}$, then either (1) $n=O(\log T)$ is sufficiently small that $\mathrm{poly}(T)$ queries to $\mathcal{U}_{n}$ actually suffice to learn $\mathcal{U}_{n}$ to inverse-polynomial precision, or else (2) $n=\omega(\log T)$ is sufficiently large that with high probability, the maximum acceptance probability of $\mathcal{V}$ (over the choice of Merlin’s witness) is close to the average maximum acceptance probability of $\mathcal{V}$ when $\mathcal{U}_{n}$ is replaced by a random set of matrices sampled from the Haar measure. We prove this as a consequence of the extremely strong concentration of measure properties exhibited by the Haar measure [Mec19].
@@
For a certain carefully-constructed language $\mathcal{C}$, this allows a $\mathsf{BQP}^{\mathcal{U},\mathcal{C}}$ machine to approximate the maximum acceptance probability of $\mathcal{V}^{\mathcal{U},\mathcal{C}}$ as follows. In case (1), the $\mathsf{BQP}^{\mathcal{U},\mathcal{C}}$ machine first queries $\mathcal{U}_{n}$ enough times to learn a unitary transformation $\widetilde{\mathcal{U}}_{n}$ that is close to $\mathcal{U}_{n}$, and then hard codes $\widetilde{\mathcal{U}}_{n}$ into a new $\mathsf{QMA}^{\mathcal{C}}$ verifier $\mathcal{W}^{\mathcal{C}}$ that simulates $\mathcal{V}$ by replacing queries to $\mathcal{U}_{n}$ with calls to $\widetilde{\mathcal{U}_{n}}$. In case (2), the $\mathsf{BQP}^{\mathcal{U},\mathcal{C}}$ machine similarly constructs a new $\mathsf{QMA}^{\mathcal{C}}$ verifier $\mathcal{W}^{\mathcal{C}}$, instead simulating $\mathcal{V}$ by replacing queries to $\mathcal{U}_{n}$ with queries to independently chosen Haar-random unitaries $\overline{\mathcal{U}}_{n}$. Thus, the problem of approximating the maximum acceptance probability of $\mathcal{V}^{\mathcal{U},\mathcal{C}}$ reduces to approximating the maximum acceptance probability of $\mathcal{W}^{\mathcal{C}}$, averaged over $\overline{\mathcal{U}}_{n}$. The language $\mathcal{C}$ is constructed in such a fashion that querying $\mathcal{C}$ on a description of $\mathcal{W}$ returns the desired approximation.
@@ 1.4 Open problems
Can we prove a similar result to Theorem 2 using a classical oracle, for either PRUs or PRSs? Attempting to resolve this question seems to run into many of the same difficulties that arise in constructing a classical oracle separation between $\mathsf{QCMA}$ and $\mathsf{QMA}$, which also remains an open problem [AK07]. For one, as pointed out in [AK07], we do not even know whether every $n$-qubit unitary transformation can be approximately implemented in $\mathrm{poly}(n)$ time relative to some classical oracle—this is sometimes known as the unitary synthesis problem [Aar16, Ros21, LMW24]. Even if one could resolve this, it is not clear whether the resulting PRUs or PRSs would be secure against adversaries with the power of $\mathsf{QMA}$. For instance, we show in Appendix A that an existing construction of PRSs, whose security is provable in the random oracle model [BS19], can be broken with an $\mathsf{NP}$ oracle. Nevertheless, recent work by Kretschmer, Qian, Sinha, and Tal [KQST23] makes progress on this question by constructing an oracle relative to which $\mathsf{P}=\mathsf{NP}$ and a weaker version of pseudorandom states (with only single-copy security) exist.
What else can be said about the hardness of learning quantum states and unitary transformations, either in the worst case or on average? A related question is to explore the hardness of problems involving quantum meta-complexity: that is, problems that themselves encode computational complexity or difficulty. Consider, for example, a version of the minimum circuit size problem ($\mathsf{MCSP}$) for quantum states: given copies of a pure quantum state $|\psi\rangle$, determine the size of the smallest quantum circuit that approximately outputs $|\psi\rangle$. If PRSs exist, then this task should be hard, but placing an upper bound on the complexity of this task might be difficult in light of our results. We view this problem as particularly intriguing because it does not appear to have an obvious classical analogue, and also because of its relevance to the wormhole growth paradox and Susskind’s Complexity=Volume conjecture in AdS/CFT [BFV20, Sus16b, Sus16a]. A number of recent breakthroughs in complexity theory have involved ideas from meta-complexity (see surveys by Allender [All17, All20] or Lu and Oliveira [LO22]), and it would be interesting to see which of these techniques could be ported to the quantum setting.
@@
1.5 Conference version
This paper improves upon the earlier conference version [Kre21a] in two major ways. First, the section on $t$-designs (Section 3) is a new addition. Second, the oracle $\mathcal{O}=(\mathcal{U},\mathcal{C})$ that we use to collapse $\mathsf{PromiseQMA}^{\mathcal{U},\mathcal{C}}$ to $\mathsf{PromiseBQP}^{\mathcal{U},\mathcal{C}}$ is different. In both versions, the quantum part $\mathcal{U}$ is the same. However, the conference version claimed that the classical oracle $\mathcal{C}$ could be any $\mathsf{PSPACE}$-complete language. By contrast, in this paper $\mathcal{C}$ is a specific recursively-constructed language. The reason for this change is an error that was discovered in the proof. We discuss this error, along with the prospects of restoring the original oracle construction, in Section 5.4.
@@ 2.1 Basic notation
2.2 Probability
We require two basic facts about probability. The first regards the optimality of the Bayes decision rule, which is a strategy for guessing a random variable $X$ from posterior information $Y$. The Bayes decision rule is to always guess the value of $X$ that maximizes the posterior probability given $Y$. The Bayes decision rule is optimal, in the sense that any other strategy that guesses $X$ using $Y$ errs at least as often as the Bayes decision rule. We only need the following special case of this fact, which also applies more generally (see [Ber13, Chapter 4.4.1] for further discussion).
Lemma 5 (Bayes decision rule).
Let $X$ be a $\{0,1\}$-valued random variable, let $Y$ be a random variable (not necessarily independent of $X$) with domain $D$, and let $f:D\to\{0,1\}$. Then:
$\Pr[f(Y)=X]\leq\Pr\left[\operatorname*{arg\,max}_{i}\Pr[X=i|Y]=X\right].$
The other fact we need is the Borel–Cantelli lemma for sequences of probabilistic events. It gives a criterion under which at most finitely many of the events occur, with probability $1$.
Lemma 6 (Borel–Cantelli [Bor09, Can17]).
Let $\{X_{n}\}_{n\in\mathbb{N}}$ be a sequence of (not necessarily independent) random variables with values in $\{0,1\}$. If
$\sum_{n=1}^{\infty}\mathop{\mathbb{E}}[X_{n}]<\infty,$
then
$\Pr\left[\sum_{n=1}^{\infty}X_{n}=\infty\right]=0.$
@@ 2.3 Quantum information
We let $\mathrm{TD}(\rho,\sigma)$ denote the trace distance between density matrices $\rho$ and $\sigma$. For a matrix $M$ we use $||M||_{F}\coloneqq\sqrt{\mathrm{Tr}\left(M^{\dagger}M\right)}$ to denote its Frobenius norm.
@@
$\left|\left|\mathcal{A}\right|\right|_{\diamond}\coloneqq\sup_{\mathrm{Tr}(\rho)=1,\rho\succeq 0}||(\mathcal{A}\otimes I)(\rho)||_{1},$
@@
$\mathrm{TD}(\mathcal{A}(\rho),\mathcal{B}(\rho))\leq\frac{1}{2}\left|\left|\mathcal{A}-\mathcal{B}\right|\right|_{\diamond}.$
@@
A consequence of 8 is the following well-known bound relating distance in diamond norm to the Frobenius norm for unitary channels. We provide a proof for completeness.
@@
Let $d$ be the distance in the complex plane between 00 and the polygon whose vertices are $\lambda_{1},\ldots,\lambda_{N}$. Then from 8 we may conclude: